OpenStudy (anonymous):

A 75.0 kg astronaut is training for accelerations that he will experience upon reentry. He is placed in a centrifuge (r = 8.0 m) and spun at a constant angular velocity of 25.3 rpm. Answer the following: a. What is the angular velocity of the centrifuge in rad/s? b. What is the linear velocity of the astronaut at the outer edge of the centrifuge? c. What is the centripetal acceleration of the astronaut at the end of the centrifuge? d. How many g’s does the astronaut experience? e. What is the centripetal force and net torque experienced by the astronaut? Give magnitudes and directions. S

OpenStudy (loser66):

knock knock

OpenStudy (loser66):

a) apply the convert unit formula \(\Large \frac{25rev}{1min}\frac{2\pi rad}{1rev}\frac{1min}{60sec}\) and you get the answer b) velocity = \(\omega r\) c)centripetal acceleration \(\alpha=\omega^2r\) d) \(\large \alpha \frac{g}{9.8}\)to get the answer